OpenStudy (anonymous):

help!?!? Find the volume of a cylinder with a height of 2 meters and a diameter that is 5 times the measure of the height. Use 3.14 for pi and round your answer to the nearest hundredth. 157 m3 339.12 m3 678.24 m3 1570 m3

OpenStudy (anonymous):

@RadEn help plz

OpenStudy (raden):

use this formula : Volume = pi * r^2 * h with r = radius and h = height

OpenStudy (anonymous):

39.26990817

OpenStudy (raden):

ok, known that h = 2 and diameter = 5 * 2 = 10 therefore its radius = d/2 = 10/2 = 5 meters now, evaluate this V = pi * r^2 * h V = 3.14 * 5^2 * 2 V = 3.14 * 25 * 2 V = ... ?

OpenStudy (anonymous):

157

OpenStudy (raden):

you are right

OpenStudy (anonymous):

thx

OpenStudy (raden):

you're welcome